OpenStudy (anonymous):

LIM ln(1/x)^x x--->0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\ln(\frac{ 1 }{ x })^{x}\]

OpenStudy (anonymous):

i changed it to xln(1/x), and i let x=1/x such that i ended up having lnx/x , when i did the whole thing, i got infinity, wolfram has 0!

OpenStudy (anonymous):

If you let \(x=\dfrac{1}{x}\), then \(\dfrac{1}{x}\to\infty\) since \(x\to0\).

OpenStudy (solomonzelman):

I am not sure about his idea, but what I thought was that\[\int\limits_{ }^{ }\frac{1}{u}~dx~=\ln(u)\]and thus,\[\frac{d}{dx}~\ln(u)~=\frac{1}{u}\]

OpenStudy (solomonzelman):

it should be 1/x in the integral.

OpenStudy (solomonzelman):

and ln(x)

OpenStudy (solomonzelman):

but I think that I am wrong .

OpenStudy (anonymous):

which means if we differentiate ln(1/x) we get x?

OpenStudy (anonymous):

x times 1/x^2

OpenStudy (anonymous):

which is i/x

OpenStudy (anonymous):

1/x I mean

OpenStudy (anonymous):

\[\lim_{x\to0}\ln\left(\frac{1}{x}\right)^x=\lim_{x\to\infty}\ln x^{1/x}=\lim_{x\to\infty}\frac{\ln x}{x}\] L'Hopital's rule from here: \[\lim_{x\to\infty}\frac{\ln x}{x}=\lim_{x\to\infty}\frac{\frac{1}{x}}{1}=0\]

OpenStudy (anonymous):

so you have to change the limit to infinity....

OpenStudy (anonymous):

That's if you use the substitution you mentioned.

OpenStudy (anonymous):

oh thanx!, now it all makes sense, thank you so much. do you have another method?

OpenStudy (anonymous):

You can still compute the limit as \(x\to0\): \[\lim_{x\to0}\ln\left(\frac{1}{x}\right)^x=\lim_{x\to0}\ln x^{-x}=-\lim_{x\to0}\frac{\ln x}{\frac{1}{x}}\] Then using L'Hopital's \[-\lim_{x\to0}\frac{\frac{1}{x}}{-\frac{1}{x^2}}=\lim_{x\to0}\frac{\frac{1}{x}}{\frac{1}{x^2}}=\lim_{x\to0}\frac{1}{\frac{1}{x}}=\lim_{x\to0}x=0\]

OpenStudy (anonymous):

thanx @SithsAndGiggles and @SolomonZelman . i am still getting used to the indeterminate form of limits